# Enhanced Mean Field Game for Interactive Decision-Making with Varied Stylish Multi-Vehicles

### v1

@@ I Introduction
In recent years, autonomous driving has advanced rapidly, with substantial progress in perception [lin2025slam2, lin2024dpl], planning [yuan2025bio], and control [tsai2024autonomous]. Nevertheless, interactive driving scenarios remain a major challenge, where safety is paramount. As illustrated in Figure1, an autonomous vehicle (AV) navigating among human-driven vehicles (HDVs) must continuously assess risks and adjust its trajectory to avoid collisions. The growing presence of AVs in heterogeneous traffic environments has further increased the complexity of interactions, requiring decision-making frameworks that can handle diverse human driving behaviors [slade2024human, wang2024multimodal]. In dense traffic, such behaviors range from aggressive lane-changing to conservative following [zhu2025fdnet, chen2024safety], creating significant variability that traditional uniform modeling approaches fail to capture [zhao2024potential].
@@
Most existing methods assume homogeneous agent behavior, neglecting the influence of behavioral diversity on traffic flow [lin2025safety, chen2024end, lin2024conflicts]. Classical game-theoretic models, though mathematically elegant, struggle with scalability in large-scale multi-agent scenarios and cannot fully represent the nuanced variations of human drivers [tan2021risk, wu2022humanlike]. Likewise, conventional control strategies [li2025adaptive, li2025efficient] often exhibit excessive conservatism in dense environments, leading to reduced efficiency and unsafe reactive decisions. Control Barrier Functions (CBFs) provide formal safety guarantees by preventing the system from entering unsafe states [lin2024enhanced, ames2019control], such as avoiding dynamic obstacles [9483029]. However, CBFs tend to be overly restrictive, forcing AVs to yield excessively and compromising efficiency. Striking a balance between safety and performance thus remains a difficult challenge in CBF-based approaches.
@@
The emergence of data-driven methodologies has offered alternative pathways for modeling complex vehicular interactions [dai2024marp]. However, these approaches encounter significant limitations in interpretability and generalization across diverse traffic configurations [hassija2024interpreting]. Learning-based systems, despite their capacity to capture intricate behavioral patterns [tian2025evaluating, paparella2024deep, xu2024towards], require extensive training datasets and often lack the theoretical foundations necessary for safety guarantees in critical scenarios [rasol2024exploring, peng2024predictive]. Furthermore, the black-box nature of these systems impedes their deployment in safety-critical applications where decision transparency is paramount.
@@
In recent years, MFG theory has emerged as a promising framework for modeling interactive driving. Unlike traditional game-theoretic approaches, MFG approximates the aggregate behavior of a large population of vehicles using a mean field term, thereby avoiding the need to model every pairwise interaction explicitly and mitigating the scalability issues inherent in conventional game formulations. However, existing MFG-based methods for autonomous driving often suffer from several limitations: they lack simulations across diverse traffic scenarios, fail to compare against other game-theoretic baselines, and remain disconnected from control-level optimization. A notable exception is [zheng2025mean], which combines MFG with MPC-QP to capture collective vehicle dynamics while applying a receding-horizon optimization scheme for real-time control. This integration addresses the conservatism of Stackelberg games and the computational burden of Nash games, providing a more flexible and scalable solution for interactive driving. Nevertheless, most prior MFG models overlook the heterogeneity of driving styles among surrounding vehicles, limiting their realism and applicability in real-world traffic. Furthermore, they fail to ensure safe interactions in densely populated environments with varied behavioral agents—an essential capability for real-world deployment.

### v2

@@ I Introduction
In recent years, autonomous driving has advanced rapidly, with substantial progress in perception [1, 2], planning [3], and control [4]. Nevertheless, interactive driving scenarios remain a major challenge, where safety is paramount. As illustrated in Figure1, an autonomous vehicle (AV) navigating among human-driven vehicles (HDVs) must continuously assess risks and adjust its trajectory to avoid collisions. The growing presence of AVs in heterogeneous traffic environments has further increased the complexity of interactions, requiring decision-making frameworks that can handle diverse human driving behaviors [5, 6]. In dense traffic, such behaviors range from aggressive lane-changing to conservative following [7, 8], creating significant variability that traditional uniform modeling approaches fail to capture [9].
@@
Most existing methods assume homogeneous agent behavior, neglecting the influence of behavioral diversity on traffic flow [10, 11, 12]. Classical game-theoretic models, though mathematically elegant, struggle with scalability in large-scale multi-agent scenarios and cannot fully represent the nuanced variations of human drivers [13, 14]. Likewise, conventional control strategies [15, 16] often exhibit excessive conservatism in dense environments, leading to reduced efficiency and unsafe reactive decisions. Control Barrier Functions (CBFs) provide formal safety guarantees by preventing the system from entering unsafe states [17, 18], such as avoiding dynamic obstacles [19]. However, CBFs tend to be overly restrictive, forcing AVs to yield excessively and compromising efficiency. Striking a balance between safety and performance thus remains a difficult challenge in CBF-based approaches.
@@
The emergence of data-driven methodologies has offered alternative pathways for modeling complex vehicular interactions [20]. However, these approaches encounter significant limitations in interpretability and generalization across diverse traffic configurations [21]. Learning-based systems, despite their capacity to capture intricate behavioral patterns [22, 23], require extensive training datasets and often lack the theoretical foundations necessary for safety guarantees in critical scenarios [24, 25]. Furthermore, the black-box nature of these systems impedes their deployment in safety-critical applications where decision transparency is paramount.
@@
In recent years, MFG theory has emerged as a promising framework for modeling interactive driving. Unlike traditional game-theoretic approaches, MFG approximates the aggregate behavior of a large population of vehicles using a mean field term, thereby avoiding the need to model every pairwise interaction explicitly and mitigating the scalability issues inherent in conventional game formulations. However, existing MFG-based methods for autonomous driving often suffer from several limitations: they lack simulations across diverse traffic scenarios, fail to compare against other game-theoretic baselines, and remain disconnected from control-level optimization. A notable exception is [26], which combines MFG with MPC-QP to capture collective vehicle dynamics while applying a receding-horizon optimization scheme for real-time control. This integration addresses the conservatism of Stackelberg games and the computational burden of Nash games, providing a more flexible and scalable solution for interactive driving. Nevertheless, most prior MFG models overlook the heterogeneity of driving styles among surrounding vehicles, limiting their realism and applicability in real-world traffic. Furthermore, they fail to ensure safe interactions in densely populated environments with varied behavioral agents—an essential capability for real-world deployment.